OpenStudy (anonymous):

Raffle tickets. One thousand raffle tickets are sold for $1 each. One prize of $800 is to be award a)Rena Condos purchases one ticket. Determine her expected value. b) Determine the fair price of a ticket.

OpenStudy (anonymous):

The probabilities are 1/1000 and 999/1000 Multiplying and add you get: 800(1/1000)+x(999/1000) = 0 800 + 999x = 0 x = -(800/999) dollars x = -0.80 dollars The price to play should be 80 cents.

OpenStudy (anonymous):

that helps

OpenStudy (anonymous):

thanks Lexussmi is that the answer!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks u so much!

OpenStudy (anonymous):

welcome